OpenStudy (anonymous):

find the curl of the gradient w=x^2 cos 3x j

OpenStudy (anonymous):

0

OpenStudy (anonymous):

curl(grad(f))=0

OpenStudy (anonymous):

but how i can solve it ?

ganeshie8 (ganeshie8):

\[ Curl(\overrightarrow{F})= \nabla \times \overrightarrow{F} = \left| \begin{array}{ccc} \hat{i} & \hat{j} & \hat{k} \\ \\ \dfrac{\partial}{\partial x} & \dfrac{\partial}{\partial y} & \dfrac{\partial }{\partial z} \\ \\ P & Q & R \end{array} \right| \]

ganeshie8 (ganeshie8):

expand and plugin the partials